OpenStudy (anonymous):

Help anyone? Candice deposited $3,587.39 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $5,000? 8 years, 0 months 16 years, 7 month 8 years, 1 month 9 years, 10 months

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

\[\Large A = P(1 + \frac{ r }{ n })^{nt}\] P = $3,587.39 A = $5,000 r = 4.2/100 = 0.042 n = 2 Plug all the numbers in and let me know what you get.

OpenStudy (anonymous):

I got 0

OpenStudy (anonymous):

?

OpenStudy (ranga):

You don't have to solve for t yet. Just plug in all the numbers, simplify and let me know what you get. Then we will do the next step and solve for t.

OpenStudy (anonymous):

I got 3739.64

OpenStudy (ranga):

You got got 3739.64 for what?

OpenStudy (anonymous):

plugging in all the numbers

OpenStudy (ranga):

Plug in the numbers into the formula: 5000 = 3587.39 * ( 1 + 0.042 / 2 )^(2t) 5000 = 3587.39 * (1.021)^(2t) Divide both sides by 3587.39 1.3938 = (1.021)^(2t) Take log on both sides log(1.3938) = log((1.021)^(2t)) log(1.3938) = 2t * log(1.021) 0.1442 = 2t * 0.0090 = 0.018t divide both sides by 0.018 8.0 = t Sp t = 8 years, 0 months.

OpenStudy (ranga):

So t = 8 years, 0 months.

OpenStudy (anonymous):

Ahhhhh thank you so much

OpenStudy (ranga):

alright.